OpenStudy (anonymous):

Jess uses her office scanner to scan pages at the rate of $0.09 per page. She decides to rent a scanner for $70 a year. The cost of scanning using the rented scanner is $0.02 per page. Part A: Write an inequality that can be used to calculate the number of pages that Jess should scan in a year so that the amount she pays for the rented scanner is less than the office scanner. Define the variable used. (5 points) Part B: How many pages should Jess scan in a year to justify renting the scanner? Show your work. (5 points)

OpenStudy (anonymous):

O = 0.09x R = 70 + 0.02x O is office cost, R is rented cost and x is the number of pages. to find the number of pages that justify, make O = R and solve for x. 0.09x = 70 + 0.02x x = 1000

OpenStudy (anonymous):

wow thxs so much \

OpenStudy (anonymous):

glad i could help ^_^

OpenStudy (anonymous):

yup >~<